8(3-m)+7=47
solve
show work

Answer by vidhyak(80)

8(3-m) + 7 = 47
8(3-m) = 47 - 7
8(3-m) = 40
3 - m = 40/8 = 5
3 - m = 5
-m = 2
m = -2